General Meeting Summary 6/30/22

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Aaron Brown, Alexandre Camsonne, Fernando Barbosa, Bogdan Wojsekhowski, Brad Sawatzky, Carlos Munoz-Camacho, Deb Biswas, Erik Wrightson, Hamlet Mkrtchyan, Julie Roche, Steven Lassiter, Marie Boer, Mark Jones, Paul King, Abishek Karki, Paulo Medeiros, Vardan Tadevosyan, Vladimir Berdnikov, Tanja Horn, Pramita Tiwari


ASSEMBLY UPDATES

  • Crystal stacking
  • viewing photos of NPS during assembly - one taken at same angle as the mechanical drawing
  • Cabling
  • presently 4 columns are fully cabled and PCBs installed
  • time for installation: 2 columns/day - 2 people can work, need to take shifts
  • cabling the remaining 26 columns would take 3 weeks total
  • cosmic paddles are installed and connectors - HV and signal cables need to be connected to the paddles
  • Ongoing:
  • cabling on top part (PS, DAQ, LED controls, LV)
  • connect HV and signal cables and PS for fans
  • DAQ/Cables/Infrastructure - 80% moved into EEL-108 and cables being bundled
  • The team: Hamlet, Vardan, Pramita, Erik
  • God progress - to date 200 cables bundles
  • Discussion about moving NPS equipment out of the HD cleanroom area with target date July 6th
  • NPS detector stays in its area
  • crates can be disposed of
  • HV cables complete in EEL126 - ready to go
  • LV cables - being put together (Fernando and team)


ASSEMBLY PLANNING

  • Finishing the cabling of 26 remaining columns
  • Orsay team could come to finish up cabling
  • in preparation need to do PMT assembly
  • Preparation of PMTs
  • put PMT together with HV divider and put sticker on it
  • Orsay team can then do the more detailed assembly
  • plan to complete PMT preparation by end of August
  • Maintenance of detector
  • anyone can participate and be training in the assembly of columns
  • contact Carlos to schedule
  • Turning on the HV and start tests
  • plan to hook up the 4 columns into the DAQ before the end of July
  • PMT checkout with scope
  • plug into DAQ
  • check out system with DAQ
  • run and read all instrument analysis and controls (chiller, temp, etc.)
  • Brad should be there for initial HV turning on
  • Check of HV cables
  • continuity check - was done for inside cables
  • recommendation: do visual inspection for outside cables
  • have some spare cables as needed


DSG UPDATES (Aaron)

  • EPICS PHOEBUS screen development - monitoring and control screens for both chiller and thermal readback
  • can be rolled out as soon as next week for checking out the readbacks before turning on the HV
  • ready to run on Linux - software environment should be compatible
  • System to run on is cdaql3 - start staging software there
  • Configuration file control
  • Ansys thermal analysis
  • investigate how heat propagates along crystals
  • Cables for scintillator paddles: all parts on site except pins; once these arrive assembly can happen
  • HV crates
  • one is still with the DSG (HallC-1); located in cleanroom in front of EEL-108
  • LED control GUI
  • setup is done, need to do programming
  • Action Items:
  • install HV controls and GUI on cdaql3 and test
  • LED control GUI programming - set up meeting with Brad
  • LED ribbon cable connectors - crimp on connectors - contact Amrit about the task (Brad)


MECHANICAL UPDATES (Paulo)

  • New beamline procurement
  • proceed with small angle beamline
  • wait with larger angle beamline procurement for simulation results


DAQ/ELECTRONICS/SOFTWARE

  • Critical path: event unblocking - ongoing discussions with David H. about possible options
  • For detector checkout all is in end
  • modified firmware etc. will be checked
  • need to go through the full chain
  • Grounding and tests of ground
  • ground connector on HV divider
  • ground on frame is available but not planning to connect it - are there any tests that could be done to verify that it is not needed before columns cabling is finished?
  • start with a clean setup for baseline
  • next check noise pickup, e.g., plug in a number of noise devices and see if pickup exists

NEXT MEETING: THURSDAY 21 JULY 2022 AT 9:00AM ET